# Periodic classical trajectories and quantum scars in many-spin systems

### v2

@@ Appendix A Numerical calculation of Lyapunov exponents
The right-hand-side of (30) normally converges to a finite value even when the true Lyapunov exponent is zero. This is because $T_{R}$, $d_{0}$, and $M$ are finite numbers. To distinguish zero Lyapunov exponents from finite but small ones, we look at the dependence of the right-hand-side of Eq.(30) on the reset time $T_{R}$. In the case of Lyapunov-unstable trajectories, the values of $\lambda$ do not depend on $T_{R}$ (as long as $|\delta\mathcal{S}(T_{R})|$ is small enough to apply the Lyapunov stability theory). When a trajectory is Lyapunov-stable, the value of $\lambda$ decreases with $T_{R}$ as
@@
Appendix B Lyapunov exponents for ergodic phase space trajectories of classical spin chains
@@
Appendix C Procedure for computing Lyapunov exponent $\lambda_{\text{S}}$ around the unstable fixed point of the separatrix
@@
As explained in Section IV.5.3, the separatrix for the class of considered periodic trajectories consists of an unstable fixed point and two branches asymptotically departing and approaching the fixed point for $t\to-\infty$ and $t\to+\infty$ respectfully (see Figs. 2 and 6). As a result, each of the two branches exhibits in the vicinity of the fixed point an atypical additional Lyapunov instability along the time flow direction. The corresponding Lyapunov exponent $\lambda_{\text{S}}$ happens to be larger than the largest Lyapunov exponents $\lambda_{\text{p}}$ for the periodic trajectories in the vicinity of the fixed point. Below we describe how $\lambda_{\text{S}}$ is computed.
We choose $J=J^{*}$ and then follow the time evolution of a phase trajectory $\mathcal{S}(t)$ starting from the initial conditions $\mathcal{S}(0)=\mathcal{S}^{*}+\delta\mathcal{S}_{0}$, where $\delta\mathcal{S}_{0}$ is a small perturbation, meaning that $|\delta\mathcal{S}_{0}|\equiv d_{0}\ll 1$. Since the fixed point $\mathcal{S}^{*}$ is itself a phase space trajectory, the growth of $|\mathcal{S}(t)-\mathcal{S}^{*}|$ should asymptotically become proportional to $e^{\lambda_{\text{S}}t}$. Indeed, we see such a behavior in Fig. 24, where the same exponential growth is exhibited at later times by $|\mathcal{S}(t)-\mathcal{S}^{*}|$ independent of the initial deviation $\delta\mathcal{S}_{0}$. The value of $\lambda_{\text{S}}$ is then extracted as the rate of that growth.
Appendix D Functional form of $\lambda_{\text{p}}(J)$ near the separatrix $J=J^{*}$
@@
Let us start by noting that the largest Lyapunov exponent for a periodic trajectory can be considered as the time-average of the instantaneous “local stretching rates” $\lambda_{\text{inst}}(t)\equiv\lim_{\Delta t\to 0}\frac{1}{\Delta t}\log\frac{|\Delta\mathcal{S}(t+\Delta t)|}{|\Delta\mathcal{S}(t)|}$ along the trajectory:
@@
where $T$ is the period of the trajectory. We further note that the separatrix has a symmetric 8-shaped form with the unstable fixed point in the middle. The spin orientation corresponding to the fixed point is to be denoted as $\mathbf{S}^{*}$. (Note: $\mathbf{S}^{*}$ is a three-dimensional vector for the one-spin representation of a periodic trajectory, while $\mathcal{S}^{*}$ is the corresponding point in the many-spin phase space.) The rotations near the separatrix cover half of that 8-shape, while the librations cover the entire 8-shape. Since the 8-shape is symmetric, it is sufficient in the case of librations to average the stretching rates in Eq.(32) only over the half of the libration period. Furthermore, since both the half of the libration trajectory and the rotation trajectory nearly coincide in the vicinity of the separatrix, the function $\lambda_{\text{p}}(J)$ tends to be largely symmetric with respect to small deviations from the separatrix $\Delta J\equiv J-J^{*}$ in agreement with the visual impression from Figs. 7(a) and with the analysis below. Yet, this analysis also reveals a subdominant asymmetry of $\lambda_{\text{p}}(J)$.
@@
The time the phase trajectory stays near the fixed point is controlled by the smallest distance $|\Delta\mathbf{S}|_{\text{min}}$ to the fixed point $\mathbf{S}^{*}$, where $\Delta\mathbf{S}(t)\equiv\mathbf{S}(t)-\mathbf{S}^{*}$. We now show that this distance has the form
@@
The time interval $T_{S}$ of the Lyapunov growth of $|\Delta\mathbf{S}(t)|$ is then followed by the time interval $T_{0}$ of large deviations from the fixed point; that interval is approximately independent of $|\Delta J|$. At the end of $T_{0}$, $|\Delta\mathbf{S}(t)|$ crosses again the threshold value $|\Delta\mathbf{S}|_{\lambda}$ and starts approaching the fixed point in the Lyapunov regime controlled by the negative Lyapunov exponent $-\lambda_{S}$. Due to the time-reversal symmetry of the Hamiltonian dynamics, it takes another time $T_{S}$ until the trajectory arrives to the point of the closest approach to the fixed point, thereby completing the period $T$. The period $T$ can now be represented as:

### v3

@@ Appendix C Numerical calculation of Lyapunov exponents
The right-hand-side (RHS) of Eq.(47) converges to that of Eq.(33) in the limit $d_{0}\to 0$, $T_{R}\to 0$ with $M=T/T_{R}$. However, in numerical simulations, $T_{R}$ and $d_{0}$ are held finite while $M$ is increasing until the RHS of Eq. (47) approaches a stationary value. Such a procedure normally results in a finite value of $\lambda$ even when the true Lyapunov exponent is zero. To distinguish zero Lyapunov exponents from finite but small ones, we look at the dependence of the right-hand-side of Eq.(47) on the reset time $T_{R}$. In the case of Lyapunov-unstable trajectories, the values of $\lambda$ do not depend on $T_{R}$ (as long as $|\delta\mathcal{S}(T_{R})|$ is small enough to apply the Lyapunov stability theory). When a trajectory is Lyapunov-stable, the value of $\lambda$ decreases with $T_{R}$ as
@@
Appendix D Lyapunov exponents for ergodic phase space trajectories of classical spin chains
@@
Appendix E Procedure for computing Lyapunov exponent $\lambda_{\text{S}}$ around the unstable fixed point of the separatrix
@@
As explained in Section IV.5.3, the separatrix for the class of considered periodic trajectories consists of an unstable fixed point and two branches asymptotically departing and approaching the fixed point for $t\to-\infty$ and $t\to+\infty$ respectfully (see Figs. 2 and 6). As a result, each of the two branches exhibits in the vicinity of the fixed point an atypical additional Lyapunov instability along the time flow direction. The corresponding Lyapunov exponent $\lambda_{\text{S}}$ happens to be larger than the largest Lyapunov exponents $\lambda_{\text{p}}$ for the periodic trajectories in the vicinity of the fixed point. Having a fixed point as a reference trajectory implies that no multiple resets of $|\delta\mathcal{S}(t)|$ envisioned by Eq.(47) are necessary to fairly sample the Lyapunov growth. The value of $\lambda_{\text{S}}$ is, instead, computed as described below.
We choose $J=J^{*}$ and then follow the time evolution of a phase trajectory $\mathcal{S}(t)$ starting from the initial conditions $\mathcal{S}(0)=\mathcal{S}^{*}+\delta\mathcal{S}_{0}$, where $\delta\mathcal{S}_{0}$ is a small deviation from the fixed point $\mathcal{S}^{*}$, meaning that $|\delta\mathcal{S}_{0}|\equiv d_{0}\ll 1$. The growth of $|\mathcal{S}(t)-\mathcal{S}^{*}|$ is expected to become asymptotically proportional to $e^{\lambda_{\text{S}}t}$, and, indeed, such a behavior can be observed in Fig. 25, where the same exponential growth is exhibited at later times by $|\mathcal{S}(t)-\mathcal{S}^{*}|$ independent of the initial deviation $\delta\mathcal{S}_{0}$. The value of $\lambda_{\text{S}}$ is then extracted as the rate of that growth.
Appendix F Functional form of $\lambda_{\text{p}}(J)$ near the separatrix $J=J^{*}$
@@
Let us start by recalling, that according to Eq.(33), the largest Lyapunov exponent for a periodic trajectory can be expressed as the time-average of the instantaneous “local stretching rates” $\lambda_{\text{inst}}(t)$ along the trajectory:
@@
where $T$ is the period of the trajectory, and $\lambda_{\text{inst}}(t)=\lim_{\Delta t\to 0}\frac{1}{\Delta t}\log\frac{|\Delta\mathcal{S}(t+\Delta t)|}{|\Delta\mathcal{S}(t)|}$, which is the same as Eq.(34) but expressed as a limit of finite differences similar to those entering Eq.(47). We further note that the separatrix has a symmetric 8-shaped form with the unstable fixed point in the middle (see Appendix A). The spin orientation corresponding to the fixed point is to be denoted as $\mathbf{S}^{*}$. (Note: $\mathbf{S}^{*}$ is a three-dimensional vector for the one-spin representation of a periodic trajectory, while $\mathcal{S}^{*}$ is the corresponding point in the many-spin phase space.) The rotations near the separatrix cover half of that 8-shape, while the librations cover the entire 8-shape. Since the 8-shape is symmetric, it is sufficient in the case of librations to average the stretching rates in Eq.(49) only over the half of the libration period. Furthermore, since both the half of the libration trajectory and the rotation trajectory nearly coincide in the vicinity of the separatrix, the function $\lambda_{\text{p}}(J)$ tends to be largely symmetric with respect to small deviations from the separatrix $\Delta J\equiv J-J^{*}$ in agreement with the visual impression from Figs. 7(a) and with the analysis below. Yet, this analysis also reveals a subdominant asymmetry of $\lambda_{\text{p}}(J)$.
@@
The time the phase trajectory stays near the fixed point is controlled by the smallest distance $|\Delta\mathbf{S}|_{\text{min}}$ to the fixed point $\mathbf{S}^{*}$, where $\Delta\mathbf{S}(t)\equiv\mathbf{S}(t)-\mathbf{S}^{*}$: the smaller this distance, the closer $|\Delta\dot{\mathbf{S}}|$ approaches zero. We now show that this distance has the form
@@
The time interval $T_{S}$ of the Lyapunov growth of $|\Delta\mathbf{S}(t)|$ is then followed by the time interval $T_{0}$ of large deviations from the fixed point; that interval is approximately independent of small $|\Delta J|$, because $T_{0}(J)\approx T_{0}(J^{*})+\left.\frac{dT_{0}}{dJ}\right|_{J=J^{*}}\Delta J\approx T_{0}(J^{*})$. At the end of $T_{0}$, $|\Delta\mathbf{S}(t)|$ crosses again the threshold value $|\Delta\mathbf{S}|_{\lambda}$ and starts approaching the fixed point in the Lyapunov regime controlled by the negative Lyapunov exponent $-\lambda_{S}$. Due to the time-reversal symmetry of the Hamiltonian dynamics, it takes another time $T_{S}$ until the trajectory arrives to the point of the closest approach to the fixed point, thereby completing the period $T$. The period $T$ can now be represented as: